# Deconstructing the Origins of Interfacial Catalysis: Why Electric Fields are Inseparable from Solvation

### v1

@@ III.1 Phenol Electric Field: Statics and Dynamics
As indicated earlier, one of the main goals of this study is to investigate with classical empirical potentials, how the electric field on the phenol moiety is tuned and affected by environmental factors. To this end, we begin by first comparing how the electric field on an O-H bond in liquid water compares to that of the hydroxyl group of the phenol. Figure 2a) illustrates the distribution of the electric field on the center of the O-H of a water molecule in the bulk. We observe that the average field is $(1.7\pm 0.4)\times 10^{10}$ V/m consistent with previous studies examining electric fields in both bulk water using different water models Reischl20022009 , as well as in water around amino acids jong2018 .
Figure 2: a) Probability density distribution of the magnitude of the electric field at the midpoint of the O-H bond of randomly selected water molecules (dotted curve) and of phenol molecules in all the three systems studied in this work (solid lines). b) Magnitude of the electric field generated by the water molecules at the midpoint of the O-H bond of the phenol molecules for the three systems studied as a function of the distance to this bond.
Moving to the electric field on the phenol, we compared the behavior of the fields and the corresponding fluctuations along the hydroxyl group for a single phenol in the bulk, at the interface and a situation consisting of 25 phenol molecules at the surface which corresponds to a surface access of 0.7 molecules/nm2 consistent with those probed in previous experiments by Tahara and co-workers. Curiously, we observe that the electric field distributions are essentially the same - the average field is $(1.2\pm 0.4)\times 10^{10}$ V/m in all three cases. Besides, the electric field of the phenol is smaller than that compared to the one experienced by a water molecule in the bulk by 30%. Thus, in this particular case, the electric fields of the phenol at the air-water interface are not particularly different from those experienced by water molecules in the bulk and furthermore, they are also not modulated or affected by the creation of phenol clusters/aggregates at the surface of water.
@@
The preceding discussion has focused on the static properties of the electric field and the magnitude of the contributions coming from different solvation layers. However, an equally important consideration in understanding the role of electric fields in chemical reactions is the timescale of field fluctuations. Specifically, if electric fields play a role in catalyzing chemical reactions, that implies that they would essentially be the slow collective variables. To this end, we examined the equilibrium time correlation functions C(t) for the electric field computed in Figure 3 resolving also the relaxation dynamics associated with the different components along the x, y and z directions.
@@ III.2 Solvation and Electric Fields
More specifically, solvation of molecules in water is typically probed by examining the exposure or proximity of water molecules to the polar groups and the orientation of water molecules. To this end, we began by looking at the first 20 nearest neighbor (NN) water molecules to the hydroxyl group of the phenol for the three different systems. Figure 4 shows two-dimensional distributions between the magnitude of the electric field projected on the OH bond of the phenol ($|proj\bar{E}|$) and the first four NNs. In all cases, we also computed the Pearson correlation coefficient which corresponds to the number in bold shown in each panel. Note that we used the projected electric field vector on the OH phenol since this group is mostly aligned with the field as shown in the SI in Figure S2.
@@
Interestingly, for all the three systems the electric field displays a rather significant anti-correlation with the distance of the nearest neighbor water molecule. In other words, the stronger the hydrogen bond when the water-phenol distance gets shorter, the larger the field strength. This observation has been fully appreciated in previous works relating electric fields to vibrational spectroscopy - large fields along stronger hydrogen bonds result in larger red-shifts of the O-H stretch frequencyeaves2005 ; cassone2019 ; Cassone2024 . Moving to the second NN, the correlation gets slightly weaker especially for phenol in the bulk (the correlation coefficient changes from -0.63 to -0.60). For phenol at the surface of water shown in the middle and bottommost panel, the correlation is also present although we begin to see some deviation from linearity. The coupling between the field and solvation appears to significantly reduced for the third NN and is basically absent once we reach the fourth NN.
@@
Figure 4: Density distributions showing the correlation between the $|proj\bar{E}|$ and the OO distance between the phenol ($O_{PHX}$) and the first four nearest water molecules ($O_{w}$) for the a. bulk, b. 1PHX and c. 25PHX. The Pearson correlation coefficient (r) is showed in each panel.
@@
Besides the radial proximity of a water molecule to the phenol, hydrogen bonding also involves specific directionality. For this reason, we examined the orientation of the water dipoles to the O-H vector of the phenol for the first 4 NNs. In Figure 5, we illustrate the density plots showing the coupling between the field and the dipole vector of the first 4 NN water molecules. Here specifically, we show two-dimensional density distributions involving the projection of both the field and dipole along the same direction (no cross-correlation was observed between variables, see Figure S3). Since the inferences are similar for the three cases, we focus on the single phenol at the interface. The analysis of the other systems is shown in the SI (Figures S3 and S4).
@@
Figure 5: Density distributions showing the correlation between each cartesian component of the projected electric field ($projE_{i}$) and the corresponding component of the dipole vector of the first four nearest water molecules ($\mu_{i}$) in the 1PHX system.
@@
Figure 5 shows that while the average field and dipole projections along all three directions are centered around zero, these two features exhibit a positive correlation along some directions of the density plot. This behavior is also observed for the 2NN but then disappears for the 3NN and 4NN. Thus, similar to the previous analysis examining the coupling between the field and the proximity of the close hydrogen-bonded waters, we observe that the first two nearest neighbors appear to be playing the most critical role. Furthermore, the fact that the density plots associated with the 1NN and 2NN are skewed symmetrically with respect to the projection of the electric field along the O-H covalent bond of the phenol moiety implies that, in a flexible (or ab initio) model, such a bond would experience field fluctuations which produce both stretching and contraction phenomena of similar magnitude. Due to the breaking of symmetry in the z-direction due to the presence of the air-water interface, the correlation plots for the z-components shown in the bottommost panel also display the same trends but are not exactly the same as those in the X and Y direction. In particular, the observed reduced spread of the density distributions shown in the bottommost panel of Figure 5 mirrors the confinement effect along the z-axis involving reduced solvation. This aspect once again magnifies the strong correlation between the concepts of solvation and electric fields.
@@ III.3 Information Imbalance: Linking Electric Fields and Solvation
Recently, our group has shown that fluctuations in aqueous systems typically occur in a high-dimensional manifold which need to be rationalized with information theoretical techniquesDonkor_JCTC_2023 ; Donkor_JPCL_2024 . An example of this is a technique known as Information ImbalanceGlielmo_PNASNexus_2022 (II) which at its core is a statistical test aimed at assessing the relative information content of different distance measures defined on the same data set. As we mention in the Methods section, the II has been extended to a differentiable formalism (DII) which allows for easily combining heterogeneous features and determine which are the most informative subsets of these features that describe a “target” data space. Referring to Eq.3, for us to determine the most informative solvation-related collection variables that describe the electric field, we consider our “target” space, $d_{B}=\{E_{x},E_{y},E_{z}\}$, and our “feature” space, $d_{A}$ consists of a large set of solvation coordinates that we describe in more detail below. We define distances in the electric field space between points $i$ and $j$ (which are two different values of the electric field vector components at two times in our trajectory) as: $d_{B}^{ij}=||E_{x}^{i}-E_{x}^{j}||+||E_{y}^{i}-E_{y}^{j}||+||E_{z}^{i}-E_{z}^{j}||$, where ||.||||.||| | . | | denotes the Euclidean norm. Distances in $d_{A}$ are also defined similarly using all the solvation coordinates.
Figure 6: Structural variables used as feature set for the information imbalance algorithm. The same features were taken for the first 20 nearest water molecules to the phenol (considered as the $O_{PHX}O_{water}$ distance).
We focused on a particular set of features which are visually illustrated in Figure 6. For the first 20 nearest water molecules to the hydroxyl group of the phenol, we extract several physical variables probing solvation. These include the distance, orientation of these waters relative to the O-H group (characterized by the dipole vector) and finally, the angle formed by these waters and the C-O bond of the phenol ($\theta_{OOC}$). This yields a total of 100 features with which we seek to find the combination that best predicts the electric field vector along the O-H. The choice of 20 water molecules covers a length-scale of 12 Å and therefore extends up to 2-3 solvation shells. Recall that the II takes on value between 0 and 1, the former implying a set of features that can perfectly predict the quantity of interest. Thus, in the context of our current work, we seek to identify which combinations of solvation features lead to the lowest II.
Figure 7: a. Information imbalance between the $|proj\bar{E}|$ and a set of structural features. b. The same information imbalance but for a substet of high and low fluctuations of the $|proj\bar{E}|$ for the 1PHX system.
@@
Previous worksDonkor_JCTC_2023 ; Donkor_JPCL_2024 ; Wild_SciRep_2024 ; Wild_NatComm_2025 ; deltattoRobustApproachInfer2025 have shown that getting values below 0.7 is usually considered a robust sign of having variables that are very informative on the target space. While it is possible to get an “ideal” II value close to 0.0 in model systems, which is the limiting case when our target space is fully described by our selected features, in real-world data sets, this is very unlikely. The presence of noise in the data, along with hidden variables that may better describe the target space but were not considered, leads to a deviation from the condition of perfectly informative features. That being said, however, as we see in Figure 7, the II reduces from a value close to 1.0 (which represents the limiting case with uninformative features), to 0.4 as we add more meaningful variables. This allows us to determine the solvation coordinates that best predict the electric field vector.
@@
One of the arguments that has been floated in the literature regarding the role of electric fields on chemical reactivity at air-water interfaces is that it is their fluctuations that need to be taken into account and that this could be a key driver in catalysisgordon2025 ; Hao2022 . To assess the role of these fluctuations, we repeat the DII analysis focusing on the data-space where the $|proj\bar{E}|$ arises exclusively from the tails of the distribution. We focused specifically on the regime that was greater than approximately ±$\pm 1.5\sigma$ of the distribution, in other words, electric fields less than $6\times 10^{9}$ and greater than $1.6\times 10^{10}$ V/m. From this constrained dataset we repeated our analysis looking for the best set of features that leads to an optimal prediction of the electric field which is shown in Figure 7 b) (the values of the II corresponding to this plot can be found in Table S2). Interestingly, we observe that using the high-field region of the data (fields greater than $1.6\times 10^{10}$ V/m), the II plateaus at a slightly lower value when using the entire 100 features compared to what is observed in the left panel of Figure 7. We thus included two more features into the analysis, which extends the set of water molecules considered beyond the 4th NN up to the 11th NN (the complete set of best features can be found in Table S3). Looking at the correlation between the $|proj\bar{E}|$ and the distance to the 4 nearest water molecules in Fig. S7 in the SI, it becomes clear that values of the electric field from the tails of the distribution yields stronger correlations with the nearest neighbor water distances. This in turn leads to a lower value of the II. On the other hand, for the lower fluctuations part of the distribution (fields less than $6\times 10^{9}$) the II based on solvation coordinates, is much higher and the correlation of the $|proj\bar{E}|$ with the first 4 nearest water molecules is less pronounced (see bottom panel of Fig. S7 in the SI). The intuition behind this is that when the phenol begins to detach from the surface, it less solvent exposed and therefore will naturally experience weaker fields.
@@ IV Discussion and Conclusions
Moving water from bulk to microdroplet conditions has been shown to significantly accelerate a broad class of chemical reactions. The microscopic origins of this effect is thought to be induced by anomalous properties of the air-water interface. However, the exact origins of this enhanced reactivity remain hotly debated and whether they actually arise from intrinsic effects due to water at the surface, remain under serious scrutiny. Among the various leading arguments that have been put forward, the presence of large electric fields and particular solvation effects within the interfacial region have been suggested to be key players in the underlying source of accelerated reactions.
@@
While the current work has focused on a model system, phenol, to study the behavior and relationships between electric fields and solvation, we speculate that these observations are rather generic and apply to most chemical systems especially those that consist of at least one polar moiety. We thus posit that the notion of the importance of electric fields in enhancing chemical reactions has perhaps been over-sensationalized in the community, a circumstance possibly rooted on the evidence that strong externally applied electric fields are capable of profoundly modifying reaction networks and free energy landscapes. In the case of hydrogen-peroxide for example, the Mishra group has very rigorously shown that its apparent enhanced yields in microdroplets has nothing to do with anything special about water at interfacesmishra2022 . Whether this extends to other reactions, however, remains an open question. At the same time, it is also worth noting that surface potentials of interfaces extracted from classical empirical models are different compared to those treated with ab initio methodskathamnn2008 ; kathmann2011 ; wang2025charge . It would be interesting in the future to use the information theoretic techniques to examine the relationships between the electric fields and solvation arising from models explicitly including electronic degrees of freedom and polarization.

### v2

@@ III.1 Phenol Electric Field: Statics and Dynamics
As indicated earlier, one of the main goals of this study is to investigate with classical empirical potentials, how the electric field on the phenol moiety is tuned and affected by environmental factors. To this end, we begin by first comparing how the electric field on an O-H bond in liquid water compares to that of the hydroxyl group of the phenol. Figure 2a) illustrates the distribution of the electric field on the center of the O-H of a water molecule in the bulk. We observe that the average field is $(1.7\pm 0.4)\times 10^{10}$ V/m consistent with previous studies examining electric fields in both bulk water using different water models Reischl et al. (2009); Corcelli et al. (2004); Auer and Skinner (2008); Ni and Skinner (2015), as well as in water around amino acids Jong and Hassanali (2018).
Figure 2: a) Probability density distribution of the magnitude of the electric field at the midpoint of the O-H bond of randomly selected water molecules (dotted curve) and of phenol molecules in all the three systems studied in this work (solid lines). b) Magnitude of the electric field generated by the water molecules at the midpoint of the O-H bond of the phenol molecules for the three systems studied as a function of the distance to this bond. As expected the decay goes as $1/r^{2}$.
Moving to the electric field on the phenol, we compared the behavior of the fields and the corresponding fluctuations along the hydroxyl group for a single phenol in the bulk, at the interface and a situation consisting of 25 phenol molecules at the surface which corresponds to a surface access of 0.7 molecules/nm2 consistent with those probed in previous experiments by Tahara and co-workers. Curiously, we observe that the electric field distributions are essentially the same - the average field is $(1.2\pm 0.4)\times 10^{10}$ V/m in all three cases. Besides, the electric field experienced by the phenol is 30% smaller than the field experienced by a water molecule. Thus, in this particular case, the electric fields of the phenol at the air-water interface are not particularly different from those experienced by water molecules in the bulk and furthermore, they are also not modulated or affected by the creation of phenol clusters/aggregates at the surface of water.
@@
The preceding discussion has focused on the static properties of the electric field and the magnitude of the contributions coming from different solvation layers. However, an equally important consideration in understanding the role of electric fields in chemical reactions is the timescale of field fluctuations. Specifically, if electric fields play a role in catalyzing chemical reactions, that implies that they would essentially be the slow collective variables. To this end, we examined the equilibrium time correlation functions C(t) for the electric field computed by equation 6 and plotted in Figure 3 resolving also the relaxation dynamics associated with the different components along the x, y and z directions.
$C(t)=\left<\hat{E}(t)\cdot\hat{E}(0)\right$> (6)
@@ III.2 Solvation and Electric Fields
More specifically, solvation of molecules in water is typically probed by examining the exposure or proximity of water molecules to the polar groups and the orientation of water molecules. To this end, we began by looking at the 20 nearest neighbor (NN) water molecules to the hydroxyl group of the phenol for the three different systems. Figure 4 shows two-dimensional distributions between the magnitude of the electric field projected on the OH bond of the phenol ($|proj\bar{E}|$) and the four NNs. In all cases, we also computed the Pearson correlation coefficient which corresponds to the number in bold shown in each panel. Note that we used the projected electric field vector on the OH phenol since this group is mostly aligned with the field as shown in the SI in Figure S2 (the distribution of orientation for the OH bond of the phenol with respect to the z-axis is shown in figure S3).
@@
Interestingly, for all the three systems the electric field displays a rather significant anti-correlation with the distance of the nearest neighbor water molecule. In other words, the stronger the hydrogen bond when the water-phenol distance gets shorter, the larger the field strength. This observation has been fully appreciated in previous works relating electric fields to vibrational spectroscopy - large fields along stronger hydrogen bonds result in larger red-shifts of the O-H stretch frequencyEaves et al. (2005); Cassone et al. (2019); Cassone and Martelli (2024). Moving to the second NN, the correlation gets slightly weaker especially for phenol in the bulk (the correlation coefficient changes from -0.63 to -0.60). For phenol at the surface of water shown in the middle and bottommost panel, the correlation is also present although we begin to see some deviation from linearity. The coupling between the field and solvation appears to be significantly reduced for the third NN and is basically absent once we reach the fourth NN.
@@
Figure 4: Density distributions showing the correlation between the $|proj\bar{E}|$ and the $O$ distance between the phenol ($O_{PHX}$) and the four nearest water molecules ($O_{w}$) for the a. bulk, b. 1PHX and c. 25PHX. The Pearson correlation coefficient (r) is showed in each panel.
@@
Besides the radial proximity of a water molecule to the phenol, hydrogen bonding also involves specific directionality. For this reason, we examined the orientation of the water dipoles to the O-H vector of the phenol for the 4 NNs. In Figure 5, we illustrate the density plots showing the coupling between the field and the dipole vector of the 4 NN water molecules. Here specifically, we show two-dimensional density distributions involving the projection of both the field and dipole along the same direction (no cross-correlation was observed between variables, see Figure S3). Since the inferences are similar for the three cases, we focus on the single phenol at the interface. The analysis of the other systems is shown in the SI (Figures S3 and S4).
@@
Figure 5: Density distributions showing the correlation between each cartesian component of the projected electric field ($projE_{i}$) and the corresponding component of the dipole vector of the four nearest water molecules ($\mu_{i}$) in the 1PHX system.
@@
Figure 5 shows that while the average field and dipole projections along all three directions are centered around zero, these two features exhibit a positive correlation along some directions of the density plot. This behavior is also observed for the 2NN but then disappears for the 3NN and 4NN. Thus, similar to the previous analysis examining the coupling between the field and the proximity of the close hydrogen-bonded waters, we observe that the two nearest neighbors appear to be playing the most critical role. Furthermore, the fact that the density plots associated with the 1NN and 2NN are skewed symmetrically with respect to the projection of the electric field along the O-H covalent bond of the phenol moiety implies that, in a flexible (or ab initio) model, such a bond would experience field fluctuations which produce both stretching and contraction phenomena of similar magnitude. Due to the breaking of symmetry in the z-direction due to the presence of the air-water interface, the correlation plots for the z-components shown in the bottommost panel also display the same trends but are not exactly the same as those in the X and Y direction. In particular, the observed reduced spread of the density distributions shown in the bottommost panel of Figure 5 mirrors the confinement effect along the z-axis involving reduced solvation. This aspect once again magnifies the strong correlation between the concepts of solvation and electric fields.
@@ III.3 Information Imbalance: Linking Electric Fields and Solvation
[…]
Figure 6: Structural variables used as feature set for the information imbalance algorithm. The same features were taken for the 20 nearest water molecules to the phenol (considered as the $O_{PHX}O_{water}$ distance).
We focused on a particular set of features which are visually illustrated in Figure 6. For the 20 nearest water molecules to the hydroxyl group of the phenol, we extract several physical variables probing solvation. These include the distance, orientation of these waters relative to the O-H group (characterized by the dipole vector) and finally, the angle formed by these waters and the C-O bond of the phenol ($\theta_{OOC}$). This yields a total of 100 features with which we seek to find the combination that best predicts the electric field vector along the O-H. The choice of 20 water molecules covers a length-scale of 12 Å and therefore extends up to 2-3 solvation shells. Recall that the II takes on values between 0 and 1, the former implying a set of features that can perfectly predict the quantity of interest. Thus, in the context of our current work, we seek to identify which combinations of solvation features lead to the lowest II.
Figure 7: a. Information imbalance between the $|proj\bar{E}|$ and a set of structural features for phenol molecules in the 3 analyzed system. b. The same information imbalance but for a substet of high and low fluctuations of the $|proj\bar{E}|$ for the 1PHX system. c. II analysis for water molecules in the bulk and at the interface.
@@
Previous worksDonkor et al. (2023, 2024); Wild et al. (2024, 2025); Del Tatto et al. (2025) have shown that getting values below 0.7 is usually considered a robust sign of having variables that are very informative on the target space. While it is possible to get an “ideal” II value close to 0.0 in model systems, which is the limiting case when our target space is fully described by our selected features, in real-world data sets, this is very unlikely. The presence of noise in the data, along with hidden variables that may better describe the target space but were not considered, leads to a deviation from the condition of perfectly informative features. That being said, however, as we see in Figure 7, the II reduces from a value close to 1.0 (which represents the limiting case with uninformative features), to 0.4 as we add more meaningful variables. This allows us to determine the solvation coordinates that best predict the electric field vector.
@@
One of the arguments that has been floated in the literature regarding the role of electric fields on chemical reactivity at air-water interfaces is that it is their fluctuations that need to be taken into account and that this could be a key driver in catalysisLaCour et al. (2025); Hao et al. (2022). To assess the role of these fluctuations, we repeat the DII analysis focusing on the data-space where the $|proj\bar{E}|$ arises exclusively from the tails of the distribution. We focused specifically on the regime that was greater than approximately $\pm 1.5\sigma$ of the distribution, in other words, electric fields less than $6\times 10^{9}$ and greater than $1.6\times 10^{10}$ V/m. From this constrained dataset we repeated our analysis looking for the best set of features that leads to an optimal prediction of the electric field which is shown in Figure 7 b) (the values of the II corresponding to this plot can be found in Table S2). Interestingly, we observe that using the high-field region of the data (fields greater than $1.6\times 10^{10}$ V/m), the II plateaus at a slightly lower value when using the entire 100 features compared to what is observed in the left panel of Figure 7. We thus included two more features into the analysis, which extends the set of water molecules considered beyond the 4th NN up to the 11th NN (the complete set of best features can be found in Table S3). Looking at the correlation between the $|proj\bar{E}|$ and the distance to the 4 nearest water molecules in Fig. S8 (upper panel) in the SI, it becomes clear that values of the electric field from the tails of the distribution yields stronger correlations with the nearest neighbor water distances. This in turn leads to a lower value of the II. On the other hand, for the lower fluctuations part of the distribution (fields less than $6\times 10^{9}$) the II based on solvation coordinates, is much higher and the correlation of the $|proj\bar{E}|$ with the 4 nearest water molecules is less pronounced (see bottom panel of Fig. S8 in the SI). The intuition behind this is that when the phenol begins to detach from the surface, it is less solvent exposed and therefore will naturally experience weaker fields.
As a final comparison, we performed the same II analysis for the electric fields acting at the midpoint of the OH bond in water molecules located in the bulk and at the interface. To this end, we calculated the Gibbs dividing surface (GDS) in the 1PHX system and, for each frame, selected one random molecule within $GDS\pm 2$ Å and one random molecule in the bulk region. For both cases, we evaluated the electric field at the midpoint of one OH bond. We then repeated the analysis carried out for phenol: projection of the field onto the OH bond, calculation of the correlation of the field with respect to OO distances (Fig. S10), and correlation with the dipole components (Fig. S11 and S12) of the four nearest neighbors using the II analysis shown in Fig. 7c. For the II analysis we employed analogous descriptors as in the phenol case, with the only modification being the replacement of the angle definition from $\theta_{OOC}$ to $\theta_{OOH}$, where H corresponds to the hydrogen atom of the target molecule not used as the reference point for the field calculation (see Fig. S9). As shown in figure 7, the II values are larger than those obtained for phenol. This difference can be attributed to the fact that the descriptors selected for the water analysis—chosen to mirror those used for phenol—do not fully capture the underlying physical phenomena in water. To validate this, we plotted the electric field magnitude against the distance to the four nearest oxygen atoms (analogous to Fig. 4), as well as the electric field components against the corresponding dipole moment components of the four nearest water molecules (Figs. 9–12, SI). These results show that, in the bulk, the correlations with distance remain strong up to the four nearest neighbors, whereas at the interface they weaken already at the fourth neighbor, similarly to the case of phenol. Regarding correlations with respect to the dipole moment, in the bulk only the first neighbor shows a significant correlation, which then decays rapidly, while at the interface correlations persist up to the third neighbor before fading. Altogether, these findings indicate that the molecular origin of the electric fields acting on phenol and on water molecules is governed by different features in each case. This suggests that, beyond the macroscopic $1/r^{2}$ decay of the field with distance, specific solvation characteristics must be considered when analyzing the effect of electric fields on a given system.
@@ IV Discussion and Conclusions
Moving water from bulk to microdroplet conditions has been shown to significantly accelerate a broad class of chemical reactions. The microscopic origins of this effect is thought to be induced by anomalous properties of the air-water interface. However, the exact origins of this enhanced reactivity remain hotly debated and whether they actually arise from intrinsic effects due to water at the surface, remains under serious scrutiny. Among the various leading arguments that have been put forward, the presence of large electric fields and particular solvation effects within the interfacial region have been suggested to be key players in the underlying source of accelerated reactions.
@@
While the current work has focused on a model system, phenol, to study the behavior and relationships between electric fields and solvation, we speculate that these observations are rather generic and apply to most chemical systems especially those that consist of at least one polar moiety. We thus posit that the notion of the importance of electric fields in enhancing chemical reactions has perhaps been over-sensationalized in the community, a circumstance possibly rooted on the evidence that strong externally applied electric fields are capable of profoundly modifying reaction networks and free energy landscapes. In the case of hydrogen-peroxide for example, the Mishra group has shown experimentally that its apparent enhanced yields in microdroplets may not be related to it occurring at the air water interface Gallo Jr. et al. (2022). Although this matter remains controversial in the literature of the area. Whether this extends to other reactions, however, remains an open question. At the same time, it is also worth noting that surface potentials of interfaces extracted from classical empirical models are different compared to those treated with ab initio methodsKathmann et al. (2008, 2011); Wang et al. (2025). It would be interesting in the future to use the information theoretic techniques to examine the relationships between the electric fields and solvation arising from models explicitly including electronic degrees of freedom and polarization.